OpenStudy (anonymous):

If (44)x = 432, what is the value of x?

OpenStudy (babynini):

You want to get x alone on one side. How can you do that? :)

OpenStudy (anonymous):

i can do that by subtracting i think @Babynini

OpenStudy (babynini):

hm no because x is not added to 44. it's multiplied.

OpenStudy (anonymous):

oh sorry i struggle a lot at math.

OpenStudy (babynini):

it's ok! me too lol. So since it's multiplied, we're going to divide.

OpenStudy (babynini):

divide both sides by 44 x=432/44

OpenStudy (babynini):

(we got rid of the 44 on the left side by dividing it by 44!)

OpenStudy (anonymous):

lol and okay what are we multiplying?

OpenStudy (anonymous):

oh okay

OpenStudy (babynini):

so what does that equal?

OpenStudy (anonymous):

i got on my calculator 9.8 idk how.

OpenStudy (babynini):

That's exactly correct :) good job!!

OpenStudy (anonymous):

oh really thanks!

OpenStudy (anonymous):

now what do i do?

OpenStudy (babynini):

so you just simplified the right side. and now your equation looks like this x=9.8 that's your answer :) (Do they ask for it in decimal or fraction form?)

OpenStudy (anonymous):

If (4^4)x = 4^32, what is the value of x? im very sorry i wrote the equation wrong.

OpenStudy (anonymous):

@Babynini

OpenStudy (babynini):

is this a different question?

OpenStudy (anonymous):

yes

OpenStudy (babynini):

ok :) first calculate 4^4

OpenStudy (babynini):

do you know how?

OpenStudy (anonymous):

no

OpenStudy (babynini):

ok. \[4^4\] = 4 * 4 * 4 *4

OpenStudy (babynini):

because it's 4, 4 times. so what does 4 * 4* 4*4 = ?

OpenStudy (anonymous):

it equals 16

OpenStudy (anonymous):

hello @Babynini

OpenStudy (babynini):

no, that's 4 *4

OpenStudy (anonymous):

oh

OpenStudy (babynini):

(sorry, working on my own math too :)) so here we have 4*4 = 16 16*4= ?

OpenStudy (anonymous):

oh im sorry if im disturbing you and is it 16X4 or 16 to the 4 power?

OpenStudy (babynini):

no you're all good 16 x 4

OpenStudy (anonymous):

its 64

OpenStudy (babynini):

ok, let me try to explain this better sweetie. \[4^4= 4x4x4x4\] so first we take 4x4 = 16 then we take that answer and multiply it by 4 16x4 = 64 then we still have to multiply it by 4 one more time because we've only done it 3 times. 64 x 4 = ?

OpenStudy (anonymous):

it equals 256

OpenStudy (babynini):

perfect, and that's your answer for 4^4 because now you've multiplied 4 four times. make sense?

OpenStudy (babynini):

(4^4)x = 4^32 now we have 256x=4^32

OpenStudy (anonymous):

thank you so much and again im sorry for disturbing you from your own math homework. have a nice day/night

OpenStudy (babynini):

wait! this stil has to be simplified more :)

OpenStudy (babynini):

want to finish solving it?

OpenStudy (anonymous):

oh sorry i didnt know and yes please

OpenStudy (babynini):

it's ok. so 256x=4^32

OpenStudy (babynini):

now doing 4^32 would be a pain without a calculator. do you have a scientific calc?

OpenStudy (anonymous):

no i do not :/

OpenStudy (babynini):

hrm ok, that's fine :)

OpenStudy (anonymous):

ok

OpenStudy (babynini):

let's see what we can do then, give me a moment :)

OpenStudy (anonymous):

ok

OpenStudy (babynini):

let's try getting x on one side by itself first then actually. so (4^4)x=4^32 we just did this, how do we get x by itself?

OpenStudy (anonymous):

dividing

OpenStudy (anonymous):

you have to divide both sides

OpenStudy (babynini):

perfect!

OpenStudy (babynini):

so we have \[x=\frac{ 4^{32} }{ 4^4 }\]

OpenStudy (anonymous):

ok

OpenStudy (babynini):

that can be simplified more but i'm not sure how :/

OpenStudy (anonymous):

oh i understand now

OpenStudy (anonymous):

i get now

OpenStudy (babynini):

ok!

OpenStudy (anonymous):

thanks for all your help and again sorry for interrupting you from your own math homework.

OpenStudy (anonymous):

bye see ya

OpenStudy (babynini):

no it's all good :) cya